OpenStudy (anonymous):

the 14 music club members each paid the same amount for dues. The club also $370 selling magazines subscriptions, they spend $704 to organize jazz festival. Now their bank account, which started at $0, is overdrawn by $40. How much did each member pay in dues?

OpenStudy (anonymous):

370-704= -334 -334-40= -374 -374/14= 26.71 i'm pretty sure that's the answer. i could be wrong though.

OpenStudy (anonymous):

thank you can u help me with another im lost